OpenStudy (anonymous):

what is the square root of 75q r cubed

OpenStudy (jdoe0001):

\(\bf \sqrt{75qr^3}\qquad {\color{blue}{ 75\implies 3\cdot 5\cdot 5\implies 3\cdot 5^2}} \\ \quad \\ \sqrt{75qr^3}\implies \sqrt{3\cdot 5^2\cdot r^2\cdot r}\)

OpenStudy (jdoe0001):

woops missed the "q" there \(\bf \sqrt{75qr^3}\qquad {\color{blue}{ 75\implies 3\cdot 5\cdot 5\implies 3\cdot 5^2}} \\ \quad \\ \sqrt{75qr^3}\implies \sqrt{3\cdot 5^2\cdot q\cdot r^2\cdot r}\)